A manufacturer of sprinkler systems designed for fire protection claims that the average activating temperature is 135 to test this claim, you randomly selected a sample of 32 systems and find evidence to reject the manufacturer's be 135.7 with a standard deviation of 3.3 at a=0.10, do you have enough evidence to reject the manufacturer's claim?


1
Expert's answer
2021-06-08T17:15:43-0400

Hypothesized Population Mean "\mu=135"

Population Standard Deviation "\sigma=3.3"

Sample Size "n=32"

Sample Mean "\bar{x}=135.7"

Significance Level "\alpha=0.10"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu=135"

"H_1: \mu\not=135"

This corresponds to two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.10," and the critical value for two-tailed test is "z_c=1.6449."

The rejection region for this two-tailed test is "R=\{z:|>1.6449\}."


The "z" - statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{135.7-135}{3.3/\sqrt{32}}\approx1.19994"

Since it is observed that "|z|=1.19994<1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "135," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for two-tailed, the significance level "\alpha=0.10, z=1.19994," is "p=2P(Z>1.19994)=0.230163," and since "p=0.230163>0.10=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "135," at the "\alpha=0.10" significance level.